Below is a list of shipping costs quoted by six online retailers for the same item. Compute the MAD for the shipping cost data. $3.00 $2.84 $2.40 $3.02 $2.80 $2.51
Answer #1

MAD is the Mean Absolute Deviation.

MAD is calculated using the below formula:

ΜAD-ΣΣa-μ

where μ can be calculated using the formula:

μ = (3+2.84+2.4+3.02+2.8+2.51)/6 = 2.761667

MAD = (|3-2.761667|+|2.84-2.761667|+|2.4-2.761667|+|3.02-2.761667|+|2.8-2.761667|+|2.51-2.761667|)/6 = 0.2044444

Xi Xi - μ |Xi - μ|
3 0.238333 0.238333
2.84 0.078333 0.078333
2.4 -0.36167 0.361667
3.02 0.258333 0.258333
2.8 0.038333 0.038333
2.51 -0.25167 0.251667
2.761667 0.204444

MAD = 0.20444444